OpenStudy (anonymous):

For x = 1/2 the second derivative of f(x) = 2x-2 is

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you mean evaluate the 2nd derivative at x = 1/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you get the 2nd der?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

1st der?

OpenStudy (anonymous):

not getting it

OpenStudy (anonymous):

not getting what?

OpenStudy (anonymous):

the first derivative

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ f \left( x+h \right) -f \left( x \right)}{ h }\] have you done these before?

OpenStudy (anonymous):

no i havent started class yet, trying to get a jump start

OpenStudy (anonymous):

you're jumping a bit far... have you looked over limits yet?

OpenStudy (anonymous):

ok thanks